how do I find the slope and y intercept of y=-1/7x-7/8
Answer by jim_thompson5910(35256)

The equation is in y = mx+b form where m = -1/7 and b = -7/8


So this means that the slope is -1/7 and the y-intercept is (0, -7/8)
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------